OpenStudy (baldymcgee6):

Inverse function of y = 2x + ln(x)

OpenStudy (baldymcgee6):

And I don't know what the W lambert function is...

OpenStudy (anonymous):

you have to switch x and y in the equation and then solve for y

OpenStudy (baldymcgee6):

y = 2x + ln(x) Interchange x and y and re-solve for y. x = 2y + ln(y) ==> e^x = e^[2y + ln(y)] ==> e^x = e^(2y)*e^[ln(y)] ==> e^x = y*e^(2y) ==> 2*e^x = 2y*e^(2y) ==> W(2*e^x) = 2y <== y = x*e^x <=> x = W(y) ==> W(2*e^x)/2 = y Therefore, the inverse is: f^-1(x) = W(2*e^x)/2

OpenStudy (baldymcgee6):

@CliffSedge

OpenStudy (anonymous):

Gimme a sec. I lost my pencil.

OpenStudy (baldymcgee6):

That is an answer off yahoo answers... Is this so called w-function needed?

OpenStudy (baldymcgee6):

lol..

OpenStudy (anonymous):

I'm not sure if it's needed, but it can be used.

OpenStudy (anonymous):